OpenStudy (anonymous):

Solve 3^4x-7 = 4^2x+3

OpenStudy (anonymous):

3^(4x-7) = 4^(2x+3) or 3^(4x) -7 = 4^(2x) +3 ?

OpenStudy (anonymous):

Please we have answered this assignement already. Do a small search you'll find neet answers log exo hey :)

OpenStudy (anonymous):

3^(4x-7) = 4^(2x+3)

OpenStudy (rulnick):

See http://openstudy.com/updates/4f47daade4b00aeb3e3f01e2 I solved it there and provided some details.

OpenStudy (rulnick):

If you have questions just let me know.

OpenStudy (anonymous):

can you please just copy and paste your answer for me on my account? i need your answer here for my school assignment [:

OpenStudy (rulnick):

x = (-6 log2-7 log3) / (4 (log2-log3)), or about 7.31.

OpenStudy (anonymous):

can you explain to me step by step on how you got your answer please [:

OpenStudy (rulnick):

Well, for that please follow the link. It's a lot to retype (and this would be the 3rd or 4th time for me!). http://openstudy.com/updates/4f47daade4b00aeb3e3f01e2

OpenStudy (anonymous):

okay thanks for your help!